#include<stdio.h>
int f(int a,int b);
int main(){
int a=8, b=4,c=2;
b = f(c,a);
printf("a=%d,b=%d\n,a,b,c);
c = f(a+3,b-2);
printf("a=%d,b=%d,c=%d\n",a,b,c);
return 0;
}
int f(int a, int b){
a=b+6
b=a-2
printf("a=%d,b=%d\n",a,b);
return a+b;
}
what is the first line of output produced?
what is the second line of output?
what is the third line if output?
what is the forth line of output?
if there is problems ill post the choices thank you
Answer #1

Solution:

(a)

Explanation:

Finding first line of output:

=>We can see that main() function is called first in the program and inside main() function 3 variable a, b and c are declared as integer type and assigned values as below-

a = 8, b = 4 and c = 2

=>Now b = f(c, a) is called so value of c will be copied to variable a declared in the function f and value of a will be copied to variable b declared in the f so the values of variables inside function f are below-

a = 2, b = 8

=>Now a = b + 6 is executed means a = 8 + 6 => a = 14

=>Now b = a-2 will be executed means b = 14 - 2 => b = 12

=>now printf("a=%d,b=%d\n",a,b); will be executed and value of variable a and b will be printed and after printing the value line will be changes as \n is there.

=>Hence output is: a=14,b=12

(b)

Explanation:

Finding second line of output:

=>Value a + b will be returned to the main() function from function f.

=>a + b = 14 + 12

=>a + b = 26

=>So b = f(c, a); will now be completed and b = 26

=>Now printf("a=%d,b=%d,c=%d\n",a,b,c); will be executed and values of variables a, b and c will be taken from main() function and then line will be changes using \n.

=>Hence output is: a=8,b=26,c=2

(c)

Explanation:

Finding third line output:

=>Now c = f(a + 3, b-2) will be called means c = f(8 + 3, 26 - 2)

=>c = f(11, 24) will be called

=>Values of variables inside function f will have values as below

a = 11, b = 24

=>a = b + 6 means a = 24 + 6 => a = 30

=>b = a - 2 means b = 30 - 2 => b = 28

=>Hence third line output is: a=30,b=28

(d)

Explanation:

>a + b = 30 + 28

=>a + b = 58

=>c = f(11, 24) is completed now.

=>c = 58

=>So final value in main() function: a = 8, b = 26 and c = 58

=>Fourth line output is: a=8,b=26,c=58

I have explained each and every part with the help of statements attached to it.